4.2 In a particular used-car yard, 25% of customers each day are female and 30% of customers have a mobile phone. If being female and having a mobile phone are independent, for an individual customer calculate the probability that:

4.2.1 They are female and have a mobile phone (3)

4.2.2 They are male and have a mobile phone (3)

4.2.3 They are female or have a mobile phone (3)
Answer #1

4.2.1 P(Female and have a mobile phone) = P(Female) * P(Have a mobile phone) = 0.25 * 0.30 = 0.075

4.2.2 P(Male and Have a mobile phone) = P(Male) * P(Have a mobile phone) = (1 - 0.25) * 0.30 = 0.225

4.2.3 P(Female or Mobile phone) = P(Female) + P(Mobile phone) - P(Both) = 0.25 + 0.30 - 0.075 = 0.475